OpenStudy (anonymous):

Given an arithmetic sequence with a7 = 19 and a12 = 54, what is a65?

OpenStudy (anonymous):

find first term and common difference at first

Parth (parthkohli):

a7, a8, a9, a10, a11, a12. Six terms, just subtract the difference.

Parth (parthkohli):

\[54 - 19 \over 6 \]Divide by 6 as there are 6 terms to find the common difference.

OpenStudy (anonymous):

make two equations from the given

Parth (parthkohli):

Oh wait. Divide by 5.

OpenStudy (anonymous):

7?

Parth (parthkohli):

Yes. That makes the common difference for ya.

OpenStudy (anonymous):

@ParthKohli that approcah is little difficult.....I mean confusing

Parth (parthkohli):

\[a_n = a_1 +(n - 1)d \]

OpenStudy (anonymous):

so i multiply 7 * 64?

Parth (parthkohli):

Yeah yeah!

Parth (parthkohli):

@sauravshakya Not to mind when the asker understands it :P

OpenStudy (anonymous):

so the answer is 448?

Parth (parthkohli):

Ya.

OpenStudy (anonymous):

thanks for the help!

Parth (parthkohli):

Wait no

Parth (parthkohli):

First you have to find a1.

Parth (parthkohli):

a7 = 19 a6 = 19 - 7 a5 = 19 - 7 - 7 a4 = 19 - 7 - 7 - 7 a3 = 19 - 7 - 7 - 7 - 7 a2 = 19 - 7 - 7 - 7 - 7 - 7 a1 = 19 - 7 - 7- 7 - 7 - 7 - 7 a1 = 19 - (7 * 6)

Parth (parthkohli):

That makes a beautiful drawing of a sail :)

OpenStudy (anonymous):

-23 and then i * 64

OpenStudy (anonymous):

haha^^

Parth (parthkohli):

\[a_{65} = -23 + 7(65 - 1) \implies -23 + (7 \times 64) \implies -23 + 448 \]

OpenStudy (angela210793):

i think it is easier if u just  so u can find a1 and d and then replace at a65=a1+(65-1)d